OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality 2x + y > 4.

OpenStudy (anonymous):

because this is strictly greater than, the boundary line be solid or dotted?

OpenStudy (anonymous):

whenever it is strictly greater than ( > ) or strictly less than ( < ) the boundary is dotted.

OpenStudy (anonymous):

if you have greater than or equal \(\large \ge \) or less than or equal \(\large \le \), then it is solid...

OpenStudy (anonymous):

so now you need to draw the dotted boundary line. you can do that easily by drawing the line through the x and y intercepts of the line: \(\large 2x+y=4 \) once you've drawn the line, you'll need to shade.. use the point (0, 0) do determine that.